OpenStudy (anonymous):

For each angle in standard position, determine one positive and one negative angle measure that is conterminal with it. In which quadrant does the terminal arm lie? (a) pi/3 (b) - 5pi /6

Directrix (directrix):

Coterminal angles are described in the attachment. http://hotmath.com/hotmath_help/topics/coterminal-angles.html

Directrix (directrix):

(a) pi/3 π/3 + 2π = (7π)/3 π/3 - 2π = - (5π)/3 The terminal arm lies in the first Quadrant. @GizmoGirl04 See what you can do with part b, okay?

OpenStudy (anonymous):

for part b, -5 pi/6 + 2pi = -3pi/6 = -1pi/2 and -5pi/6 - 2pi = -7pi/6 is this correct? i can't figure out which quadrant it lies in.

Directrix (directrix):

Every time you add or subtract 2π radians (360 degrees) you end up where you started.

Directrix (directrix):

>>> for part b, -5 pi/6 + 2pi = -3pi/6 = -1pi/2 Not correct. -5 pi/6 + 2pi = -5pi/6 + 12pi/6 = 7pi/6 --> This is in the same quadrant as -5 pi /6 ------------- Okay, now you do the -5 pi/6 - 2pi again. @GizmoGirl04